# A Two-Stage Fourth-Order Implicit Scheme for Stiff problems

### v1

@@ 3 The Stability Analysis of the Implicit TSFO Scheme
Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable. Achieving this requires not only need condition (3.9) above but also need the following condition:
@@
The theoretical foundation for verifying A-stability is based on the maximum modulus principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$. Therefore, to achieve the A-stability of the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
@@
Condition 2: The $L_{\infty}$-norm of $G$ on the imaginary axis is bounded by 1, i.e.,
@@
Condition 3: $G$ is bounded at infinity with magnitude not exceeding 1, i.e.,
@@
The first requirement for A-stability is that the amplification factor $G(z)$ must be analytic in the closed left half-plane $\mathbb{C}_{-}$. This is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane. The denominator of $G(z)$ is given by:
@@
i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in $\textnormal{Re}(z)\leq 0$.
@@
Substitute $z=iy$ and (3.29) into $G(z)$, we obtain
@@
$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right],$
where $C\in[0.018824,0.045589]$, with $C=0.018924$ being the optimal value.
4 The Implementation of the Implicit TSFO Scheme
To achieve accelerated convergence, the implicit scheme can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [34].
@@ 4 The Implementation of the Implicit TSFO Scheme
$\quad-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
@@ 5 Numerical tests
This section will validate the convergence order, stability, and competitiveness of the proposed two-stage fourth-order implicit (e.g., Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems, with its fourth-order accuracy being verified through the extrapolation algorithm below,
$\text{Order}=\frac{\text{ln}[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})]}{\text{ln}(\Delta t_{1}/\Delta t_{2})},$ (5.1)
especially, when $\Delta t_{1}=\Delta t$, $\Delta t_{2}=\Delta t/2$, we have
$\text{Order}_{L^{2}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)]}{\text{ln}2},$ (5.2)
$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\text{ln}[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)]}{\text{ln}2},$ (5.3)
@@
$\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{2}},\quad\text{Error}_{L^{2}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{2}},$ (5.4)
$\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{e}\lVert_{L^{\infty}},\quad\text{Error}_{L^{\infty}}(\Delta t)=\lVert{\bf u}_{n}(\Delta t)-{\bf u}_{c}\lVert_{L^{\infty}},$ (5.5)
${\bf u}_{n}(\Delta t)$ is the numerical solution, ${\bf u}_{e}$ is the exact solution, and ${\bf u}_{c}$ is the reference solution generated by the classical fourth-order explicit Runge-Kutta method with a highly refined time step of $1.0\times 10^{-6}$. Additionally, a direct comparison is conducted between the proposed implicit TSFO scheme and the classical fourth-order implicit Runge-Kutta (e.g., Implicit RK4) method, focusing on stable time step size and convergence error, to verify the superiority of the former.
@@ 5.1 Linear System with Widely Separated Eigenvalues
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 1: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 2: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Linear Systems with Widely Separated Eigenvalues".
@@ 5.2 Robertson’s Chemical Kinetics Problem
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 3: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 4: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Robertson’s Chemical Kinetics Problem".
@@ 5.3 Ozone Decomposition Reaction Problem
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 5: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 6: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=1.0$.
@@
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 7: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 8: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=10.0$.
@@
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 9: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 10: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Ozone Decomposition Reaction Problem" with $t_{\text{end}}=321.8122$.
@@ 5.4 Van der Pol Oscillator Problem
$\Delta t$
${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 11: $L^{2}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
@@
$\Delta t$
${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 12: $L^{\infty}$-Norm Error and Convergence Order Analysis of Implicit TSFO and Implicit RK4 Methods for "Van der Pol Oscillator Problem".
Numerical results confirm that the proposed two-step fourth-order implicit (Implicit TSFO) scheme effectively solves stiff problems using large time steps without compromising fourth-order accuracy. When compared to the classical implicit Runge-Kutta (Implicit RK4) method, our scheme reduces the numerical error by about an order of magnitude under identical time steps and offers superior stability that enables even larger step sizes.

### v2

@@ 3 Stability Analysis for the Implicit TSFO Scheme
Our objective in the following is to determine the range of $C$ for which the Implicit TSFO scheme (2.22)-(2.23) is A-stable. Achieving this requires not only need condition (3.9) but also need the following condition:
@@
The theoretical foundation for verifying the condition (3.12) is based on the Maximum Modulus Principle: If $G(z)$ is analytic in the closed left half-plane $\text{Re}(z)\leq 0$, then the maximum of $\lvert G(z)\lvert$ occurs on the boundary, specifically on: the imaginary axis $z=iy,\;y\in\mathbb{R}$ and at infinity $z\to\infty$. Therefore, to achieve the condition (3.12) for the Implicit TSFO scheme (2.22)-(2.23), we need to examine the following three conditions:
@@
Condition 2: The $L_{\infty}$-norm of $G(z)$ on the imaginary axis is bounded by 1, i.e.,
@@
Condition 3: $G(z)$ is bounded at infinity with magnitude not exceeding 1, i.e.,
@@
Condition 1 requires the amplification factor $G(z)$ to be analytic in the closed left half-plane $\mathbb{C}_{-}$, which is equivalent to requiring that all poles of $G(z)$ lie strictly in the right half-plane. The denominator of $G(z)$ is given by:
@@
i.e., all roots of the equation (3.21) lie in the right half-plane, in another words, the amplification factor $G(z)$ is analytic in the closed left half-plane.
@@
Substitute $z=iy$ and (3.29) into (3.11), we obtain
@@
$+C\Delta t^{2}\left[\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+\frac{1}{2}})-\frac{3}{2}\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n+1})\right].$
The effective range for parameter $C$ is $[0.018824,0.045589]$, within which $C=0.018924$ serves as a favorable choice.
4 Newton’s Iteration for the Implicit TSFO Scheme
To achieve accelerated convergence, the implicit TSFO scheme (3.35)-(3.36) can be realized through Newton’s iterative method, where the initial guess for each stage uses the result of the two-stage fourth-order explicit scheme [1].
@@ 4 Newton’s Iteration for the Implicit TSFO Scheme
$\quad\quad\;\;\;-{\frac{\Delta t^{2}}{48}}\left\{\frac{\partial}{\partial t}\mathcal{L}\left[({\bf u}^{n+\frac{1}{2}})^{(k)}\right]-\frac{\partial}{\partial t}\mathcal{L}({\bf u}^{n})\right\}-({\bf u}^{n+\frac{1}{2}})^{(k)}.$
@@ 5 Numerical Tests
This section will validate the convergence order, stability, and computational efficiency of the proposed implicit two-stage fourth-order (Implicit TSFO) scheme (3.35)-(3.36) by applying it to several classic stiff problems. Its fourth-order accuracy is verified using the following extrapolation algorithm:
$\text{Order}=\frac{\ln\!\bigl[\text{Error}(\Delta t_{1})/\text{Error}(\Delta t_{2})\bigr]}{\ln\!\bigl(\Delta t_{1}/\Delta t_{2}\bigr)},$ (5.1)
in particular, when $\Delta t_{1}=\Delta t$ and $\Delta t_{2}=\Delta t/2$, we have
$\text{Order}_{L^{2}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{2}}(\Delta t)/\text{Error}_{L^{2}}(\Delta t/2)\bigr]}{\ln 2},$ (5.2)
$\text{Order}_{L^{\infty}}(\Delta t)=\frac{\ln\!\bigl[\text{Error}_{L^{\infty}}(\Delta t)/\text{Error}_{L^{\infty}}(\Delta t/2)\bigr]}{\ln 2},$
@@
$\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{2}},$ $\text{Error}_{L^{2}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{2}},$ (5.3)
$\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{e}\bigr\|_{L^{\infty}},$ $\text{Error}_{L^{\infty}}(\Delta t)$ $=\bigl\|\mathbf{u}_{n}(\Delta t)-\mathbf{u}_{c}\bigr\|_{L^{\infty}}.$
Here, $\mathbf{u}_{n}(\Delta t)$ is the numerical solution, $\mathbf{u}_{e}$ is the exact solution, and $\mathbf{u}_{c}$ is a reference solution generated by the classical fourth-order explicit Runge–Kutta method using a highly refined time step of $1.0\times 10^{-6}$. Furthermore, to demonstrate the superiority of the proposed implicit TSFO scheme, a direct comparison is conducted with the classical fourth-order implicit Runge–Kutta method (e.g., Implicit RK4), focusing on the allowable stable time step size and the corresponding convergence error.
@@ 5.1 Linear System with Widely Separated Eigenvalues
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 1: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 2: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Linear Systems with Widely Separated Eigenvalues
@@ 5.2 Robertson’s Chemical Kinetics Problem
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 3: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 4: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Robertson’s Chemical Kinetics Problem
@@ 5.3 Ozone Decomposition Reaction Problem
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 5: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 6: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=1.0$
@@
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 7: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 8: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=10.0$
@@
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 9: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 10: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Ozone Decomposition Reaction Problem at $t_{\text{end}}=321.8122$
@@ 5.4 Van der Pol Oscillator Problem
$\Delta t$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$ ${\text{Error}}_{L^{2}}(\Delta t)$ ${\text{Order}}_{L^{2}}(\Delta t)$
@@
Table 11: A Comparison of Implicit TSFO and RK4 Methods: $L^{2}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
@@
$\Delta t$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$ ${\text{Error}}_{L^{\infty}}(\Delta t)$ ${\text{Order}}_{L^{\infty}}(\Delta t)$
@@
Table 12: A Comparison of Implicit TSFO and RK4 Methods: $L^{\infty}$-Norm Error and Convergence Order for Van der Pol Oscillator Problem
The numerical experiments confirm that the proposed implicit TSFO time discretization scheme successfully achieves fourth-order temporal accuracy in only two computational stages. The results also demonstrate the robust stability of the proposed algorithm when solving stiff problems with relatively large time steps, thereby validating its effectiveness for stiff systems. In comparison to the classical fourth-order implicit Runge-Kutta method, the present scheme produces smaller errors under identical time step sizes, typically reducing the error by an order of magnitude (reference solutions are computed using the explicit fourth-order Runge-Kutta method). Moreover, it maintains stability under even larger time steps, highlighting its superior numerical stability. The source code for all test cases presented is publicly available in the accompanying GitHub repository at: https://github.com/ZhixinHuo/Implicit-TSFO.